# Aerodynamic and structural airfoil shape optimisation via Transfer Learning-enhanced Deep Reinforcement Learning

### v1

@@ 2.2.4 RL specs IV: the agent
To train an agent with reinforcement learning, one needs to choose the specific DRL algorithm. This choice may depend, among other things, on whether the action space is discrete or continuous. For continuous actions, there are a variety of algorithms, including A2C [39], PPO [40], SAC [41] or TD3 [42]. After a preliminary analysis, we decided to use PPO (Proximal Policy Optimisation), as it shows better convergence properties than A2C and TD3 and SAC have difficulties converging. PPO is a popular policy gradient algorithm that, for instance, is behind the post-training with RLHF of popular Large Language Models, underscoring its versatility. In our implementation, PPO employs neural networks to model the policy and value functions. Specifically, PPO uses an actor-critic framework where the policy network (actor) selects the actions, and the value network (critic) evaluates the expected returns of those actions. Figure 3 shows a schematic of the neural networks comprising PPO. To train the agent, the Python library Stable Baselines 3 [43] has been used. This is a library that implements different reinforcement learning algorithms and is easy to integrate into Gymnasium environments.
2.3 Transfer learning
@@
Finally, we need to explicitly discuss the transfer learning technique. We start by recalling that this is a way of accelerating the training process [28], as it has been observed in RL frameworks that an agent initially trained with a (fast) low-cost model over a substantial number of iterations and subsequently trained with a high-cost model over a limited number of iterations achieves a comparable performance with respect to training solely with a high-cost model over a longer period. Transfer learning thus reduces the overall computational cost and time required for the training process, something particularly welcome in high-cost applications such as aerodynamic shape optimisation. Here, as previously discussed, we use the surrogate model NeuralFoil as the low-cost model, and XFoil as the more accurate yet more computationally expensive solver. To provide a more effective evaluation of the efficacy of Transfer Learning we use the smallest model available of Neuralfoil, which has a clear speed advantage and a clearly lower accuracy compared to XFoil.
@@
To clarify, by ‘freezing of weights’ during fine-tuning we mean that they will not be changed. Fig. 3 illustrates how the weight transfer is performed for the specific scenario raised on strategy . Finally, note that after pre-training the first agent, an entropy term is added to the PPO loss function, which encourages the new model to explore the state space. This makes it easier for this new agent to move into other areas of the state space. The PPO configurations for each case are described in Appendix B.3.
3 Results
Results cover a variety of analysis of the DRL agents, see Appendices B.1 and B for an in-depth study of the hyperparameter selection and the final configurations of the different PPO hyperparameters.
We initially compare the performance and training’s computational efficiency of the DRL agent (trained for the single-objective, aerodynamic optimisation and trained on multi-objective, aerodynamic/structural optimisation) with respect to a traditional gradient-free optimisation methodology, namely Particle Swarm Optimisation (PSO). Then we perform a more detailed analysis of the performance of the DRL agent in the single vs multi-objective scenario. Finally, we assess the performance and efficiency gain when the DRL agent is enhanced with different possible Transfer Learning strategies.
@@ 3 Results
Once the DRL agent is trained, we evaluate its performance by running a whole episode on the set of airfoils in the Aerosandbox [44] Python library, which is essentially a superset of the UIUC Airfoil dataset. This dataset includes a broad variety of airfoils, which allows for a comprehensive evaluation of the agent’s ability to generalise across different shapes and aerodynamic characteristics.
@@
Aditionally, in the multi-objective case we monitor the deviation of the airfoil’s maximum thickness with the percentage
@@
3.1 Deep Reinforcement Learning vs Particle Swarm Optimisation
This section assesses the proposed DRL methodology and compares it (both in the single and multi-objective versions) with a traditional gradient-free method: Particle Swarm Optimisation (PSO) [45].
PSO is a population-based optimisation algorithm inspired by the social behavior of bird flocks searching for corn. It is widely used due to its simplicity, ability to handle non-differentiable objective functions, and efficiency in exploring high-dimensional search spaces. PSO has been successfully applied in various domains, including image and video analysis applications, engineering design and neural networks, and fluid dynamics optimisation. Furthermore, PSO has proven effective in airfoil optimisation [13, 14, 15].
Here, we use Xoptfoil2 [46] to perform PSO optimisation, since this tool is well designed to compare PSO and DRL as it actually makes calls to XFoil underneath.
The chosen PSO parameters are the default ones that this library provides. The results obtained using this tool are systematically compared to those given by the DRL agent trained for 81920 steps using XFoil.
@@ 3.1 Deep Reinforcement Learning vs Particle Swarm Optimisation
We start with the single-objective optimisation case, where the DRL agent is trained to solely improve the aerodynamic efficiency of the airfoil. For illustration, the first 20 steps of the DRL policy, as applied to the initial airfoil NACA0012, is shown in Fig. 4a, whereas the airfoil reaching the highest lift-to-drag ratio is depicted in Fig. 4c (center). We clearly see that the DRL agent’s policy is able to substantially and monotonically increase the aerodynamic efficiency of the airfoil.
@@
We have made available the whole optimisation trajectory for all 20 airfoils as videos: see here222https://fermingut.github.io/CETACEO_UPM/rl_optim.html and here333https://fermingut.github.io/CETACEO_UPM/pso_optim.html for the DRL and PSO trajectories, respectively.
@@
| PSO (single-objective) | 1566 | 105 50 | 198(76) |
| DRL (single-objective) | 1456 | 141 48 | 243(35) |
In a second step, we then focus on the multi-objective case. For illustration, in Fig. 4b we show the performance of a DRL agent’s policy which was trained on multi-objective, aerodynamic/structural optimisation with (see also Fig. 4c (right)). It is clear that the incorporation of the structural regularisation in the reward function has an impact: the lift-to-drag ratio is still increased (albeit reaching smaller values than in the single-objective case), but now the maximum thickness of the shape seems to be maintained.
Interestingly, within Xoptfoil2 it is possible to define geometrical constraints –such as e.g. a target maximum thickness– and therefore a fair comparison between DRL and PSO within the multi-objective case discussed in Sec. 3.2 is possible too. For illustration, in this comparison we set (that we will later show to achieve a good balance between aerodynamic optimisation and maximum thickness conservation).
@@
(videos of DRL trajectories are available for RL444https://fermingut.github.io/CETACEO_UPM/rl_optim_preserving_thickness.html and PSO555https://fermingut.github.io/CETACEO_UPM/pso_optim_preserving_thickness.html),
whereas summary statistics of the evaluation over all the converged airfoils of the UIUC airfoil dataset are reported in Table 3. While DRL remains clearly superior in terms of optimising aerodynamic efficiency, PSO seems to make a better job at preserving the maximum thickness for .
@@
| PSO (multi-objective) | 1566 | 64 36 | 150(61) | 0.67 2.21 |
| DRL (multi-objective) | 1444 | 89 37 | 179(46) | 11.5 9.4 |
Finally, we provide a comparison of optimisation time per airfoil (averaged over the whole UIUC evaluation set) for DRL and PSO, as performed on an Intel(R) Xeon(R) Gold 6248R CPU @ 3.00GHz+, with one process and one CPU thread. For concreteness, we focus on the single-objective case (where the maximum thickness is not necessarily preserved, e.g. ), although note results are qualitatively similar in the multi-objective case. The DRL agent takes about 23 seconds to optimise all 1566 airfoils, i.e. an average of 0.0147 s/airfoil. Comparatively, PSO was notably slower; actually, too slow to directly perform the optimisation in a single core, so MPI was used to accelerate the evaluation (and then the results per process were inferred). The CPU used has 48 cores, 47 used by worker processes plus the one by master process, and with this setup, optimising all the evaluation set took 22 hours and 26 minutes. Assuming linear acceleration, this means seconds on average per airfoil (using a single process). Accordingly, PSO takes about times more time than DRL to optimise each airfoil using a single process. This highlights the drastic efficiency advantage of DRL, making PSO impractically slow by comparison.
3.2 Hyperparameter exploration in the multi-objective case
Here we provide additional insights on the performance of the DRL agent trained on the multi-ojective case (i.e. in Eq. 4). We systematically vary –where larger values of provide more weight to the structural optimisation over the aerodynamic one), and assess (i) the number of training steps needed to reach training convergence, (ii) the lift-to-drag ratio performance metrics improvement and best evaluated on all the airfoils present in the Aerosandbox library and (iii) the variation in the maximum thickness, as a function of the regulariser’s hyperparameter . Results (for the converged airfoils) are depicted in Table 4. It can be observed that the number of necessary training steps varies (typically increases) with . At the same time, results confirm the evidence illustrated in Fig 4 and show that the DRL agent (i) is able to improve the aerodynamic efficiency while (ii) also maintaining the structural integrity of the airfoil –i.e. certifying the multi-objective capability–. Finally, results also certify that forcing the maximum thickness to be preserved comes at the cost of a smaller improvement of the aerodynamic efficiency.
@@ 3.2 Hyperparameter exploration in the multi-objective case
3.3 Enhancing the DRL-based airfoil optimisation with Transfer Learning
To round off, here we compare the training efficiency and evaluate the performance of the trained DRL agent when training is done using a single aerodynamic solver (XFoil alone or, for completeness, NeuralFoil alone) with respect to the case when the agent is pre-trained with NeuralFoil and fine-tuned with XFoil). For the sake of exposition, we separate the single-objective ( in Eq. 4) and the multi-objective () cases.
Single-objective case. Results for all DRL agents in the single-objective case (with and without transfer learning) are summarised in Table 5, reporting (i) the total number of training steps, (ii) the total solver time, (iii) the number of airfoils evaluated, and the metrics (iv) pointwise statistics of best and (v) improvement over all the airfoils in the evaluation set. Pointwise statistics of best reported in the table are given in terms of median (IQR) as the distributions are not Gaussian (Shapiro-Wilk test) and highly skewed, whereas in the case of improvement we report mean std as the distributions are fairly more symmetrical. Note that to compute the total solver time we simply multiply the number of training steps by the execution time per call to the particular aerodynamic solver, which according to Neuralfoil’s GitHub amounts to ms for a call to XFoil and ms for a call to the smallest version of NeuralFoil, which is the one we actually used. Although the total solver time is not equivalent to the total training time, it is nonetheless an objective way of comparing the computational effort of the agent using vs. not using transfer learning. As previously mentioned, the number of airfoils evaluated slightly varies across agents as XFoil, used universally for evaluation, sometimes has convergence issues, and we had to discard those airfoils for which XFoil did not converge.
| DRL Agent | Training steps | Total solver time (s) | Airfoils evaluated | Improvement | Best |
| Fully trained with XFoil (no transfer learning) | 81920 | 5980 | 1982 | ||
| Pre-training (Neuralfoil) | 26312 | 105 | 1970 | ||
| Transfer Learning (fine-tuning with strategy ) | 10240 | 748 | 1983 | ||
| Transfer Learning (fine-tuning with strategy ) | 10240 | 748 | 1977 |
First, out of the four transfer learning strategies depicted in Sec. 2.3, we found that strategy yielded the best improvement of training time, followed by . Strategies and did not actually yield any notable improvement and are thus not reported in the table. This suggests that the information present in the final layers of the neural network is very influential to the agent’s behaviour.
@@ 3.3 Enhancing the DRL-based airfoil optimisation with Transfer Learning
Third, not only the number of steps, but also the total solver time during training ( seconds vs. seconds) is notably shorter for the agents trained using transfer learning. On this respect, we have also compared the total training time with and without transfer learning. If we define the Time Reduction () as
we find , reaching higher reductions than in other settings [30].
@@
Multi-objective case. Results of the DRL agent (with TL) trained on a combined aerodynamic/structural optimisation are depicted in Table 6. Comparing these with the equivalent results without TL reported in Table 4, we find that (i) the procedure involving TL substantially reduces the number of training steps as well as the training time, reaching Time Reduction as large as ; (ii) fixing the structural hyperparameter , the improvement in aerodynamic efficiency is only slightly lower for every value of (around a 10%) in the TL-enhanced DRL agent than the one without TL, and (iii) the preservation of the maximum thickness is quite similar in both techniques. In summary, TL-enhancement also works in the multi-objective case.

### v2

@@ II.2.4 RL specs IV: the agent
To train an agent with reinforcement learning, one needs to choose the specific DRL algorithm. This choice may depend, among other things, on whether the action space is discrete or continuous. For continuous actions, there are a variety of algorithms, including A2C Mnih et al. (2016), PPO Schulman et al. (2017), SAC Haarnoja et al. (2018) or TD3 Fujimoto, Hoof, and Meger (2018). After running some preliminary analysis, we decided to use PPO (Proximal Policy Optimisation), as it shows better convergence properties than A2C. Moreover, we discarded the use of TD3 and SAC since they had difficulties converging. PPO is a popular policy gradient algorithm that, for instance, is behind the post-training with RLHF of popular Large Language Models, underscoring its versatility. In our implementation, PPO employs neural networks to model the policy and value functions. Specifically, PPO uses an actor-critic framework where the policy network (actor) selects the actions, and the value network (critic) evaluates the expected returns of those actions. Figure 3 shows a schematic of the neural networks comprising PPO. To train the agent, the Python library Stable Baselines 3 Raffin et al. (2021) has been used. This is a library that implements different reinforcement learning algorithms and is easy to integrate into Gymnasium environments.
II.3 Transfer learning
@@
Finally, we need to explicitly discuss the transfer learning technique. We start by recalling that this is a way of accelerating the training process Yan et al. (2019), as it has been observed in RL frameworks that an agent initially trained with a (fast) low-cost model over a substantial number of iterations and subsequently trained with a high-cost model over a limited number of iterations achieves a comparable performance with respect to training solely with a high-cost model over a longer period. Transfer learning thus reduces the overall computational cost and time required for the training process, something particularly welcome in high-cost applications such as aerodynamic shape optimisation. Here, as previously discussed, we use the surrogate model NeuralFoil as the low-cost model, and XFoil as the more accurate yet more computationally expensive solver. To provide a more effective evaluation of the efficacy of Transfer Learning we use the smallest model available of NeuralFoil, which has a clear speed advantage and a clearly lower accuracy compared to XFoil.
@@
To clarify, by ‘freezing of weights’ during fine-tuning we mean that they will not be changed. Fig. 3 illustrates how the weight transfer is performed for the specific scenario raised on strategy . Finally, note that after pre-training the first agent, an entropy term is added to the PPO loss function, which encourages the new model to explore the state space. This makes it easier for this new agent to move into other areas of the state space. The PPO configurations for each case are described in Appendix B.
III Results
Results cover a variety of analysis of the DRL agents, see Appendix B for an in-depth study of the hyperparameter selection and the final configurations of the different PPO hyperparameters.
We initially compare the performance and training’s computational efficiency of the DRL agent (trained for the purely aerodynamic optimisation and then also trained on hybrid aerodynamic/structural optimisation) with respect to a traditional gradient-free optimisation methodology, namely Particle Swarm Optimisation (PSO). Then we perform a more detailed analysis of the performance of the DRL agent in the single vs multi-objective scenario. Finally, we assess the performance and efficiency gain when the DRL agent is enhanced with different possible Transfer Learning strategies.
@@ III Results
Once the DRL agent is trained, we evaluate its performance by running a whole episode on the set of airfoils in the Aerosandbox Sharpe (2021) Python library, which is essentially a superset of the UIUC Airfoil dataset. This dataset includes a broad variety of airfoils, which allows for a comprehensive evaluation of the agent’s ability to generalise across different shapes and aerodynamic characteristics.
@@
Aditionally, in the hybrid optimisation case we monitor the deviation of the airfoil’s maximum thickness with the percentage
@@
III.1 Deep Reinforcement Learning vs Particle Swarm Optimisation
This section assesses the proposed DRL methodology and compares it (both for purely aerodynamic and hybrid optimisation) with a traditional gradient-free method: Particle Swarm Optimisation (PSO) Kennedy and Eberhart (1995).
PSO is a population-based optimisation algorithm inspired by the social behaviour of bird flocks searching for corn. It is widely used due to its simplicity, ability to handle non-differentiable objective functions, and efficiency in exploring high-dimensional search spaces. PSO has been successfully applied in various domains, including image and video analysis applications, engineering design and neural networks, and fluid dynamics optimisation. Furthermore, PSO has proven effective in airfoil optimisation Wang, Zhang, and Chen (2011); Fourie and Groenwold (2002); Nejat, Mirzabeygi, and Panahi (2014) and seems to better suited than other standard gradient-free choices such as Genetic Algorithms for airfoil optimisation Mukesh, Lingadurai, and Elamvaluthi (2012); Hoyos et al. (2021).
Here, we use Xoptfoil2 Jochen Guenzel (2025) to perform PSO optimisation, since this tool is well designed to compare PSO and DRL as it actually makes calls to XFoil underneath.
The chosen PSO parameters are the default ones that this library provides. The results obtained using this tool are systematically compared to those given by the DRL agent trained for 81920 steps using XFoil. Since the main point of this paper is DRL, we do not conduct a thorogh hyperparameter optimisation for PSO, so we cannot guarantee that its results may slightly differ after thorough hypertuning.
@@ III.1 Deep Reinforcement Learning vs Particle Swarm Optimisation
We start with the purely optimisation case, where the DRL agent is trained to solely improve the aerodynamic efficiency of the airfoil. For illustration, the first 20 steps of the DRL policy, as applied to the initial airfoil NACA0012, is shown in Fig. 4a, whereas the airfoil reaching the highest lift-to-drag ratio is depicted in Fig. 4c (center). We clearly see that the DRL agent’s policy is able to substantially and monotonically increase the aerodynamic efficiency of the airfoil.
@@
| PSO (pure aerodynamic) | 1566 | 105 50 | 198(76) |
| DRL (pure aerodynamic) | 1456 | 141 48 | 243(35) |
In a second step, we then focus on the hybrid optimisation case. For illustration, in Fig. 4b we show the performance of a DRL agent’s policy which was trained on hybrid aerodynamic/structural optimisation with (see also Fig. 4c (right)). The choice of is arbitrary, it is not aimed at signalling any Pareto optimality but rather is selected for illustration purposes. The interested practitioner will actually need to select the value of depending on the specific priorities of their application.
It is clear that the incorporation of the structural regularisation in the reward function has an impact: the lift-to-drag ratio is still increased (albeit reaching smaller values than in the purely aerodynamic case), but now the maximum thickness of the shape seems to be maintained.
Interestingly, within Xoptfoil2 it is possible to define geometrical constraints –such as e.g. a target maximum thickness– and therefore a fair comparison between DRL and PSO within the hybrid optimisation case discussed in Sec. III.2 is possible too. For illustration, in this comparison we set (that we will later show to achieve a good balance between aerodynamic optimisation and maximum thickness conservation).
@@
whereas summary statistics of the evaluation over all the converged airfoils of the UIUC airfoil dataset are reported in Table 3. It is evident that DRL exhibits superior performance in terms of aerodynamic efficiency optimisation. However, PSO demonstrates a more effective capability in preserving the maximum thickness. It’s possible that this might just be flagging the capacity of DRL to explore more design space than PSO, that in that sense only optimises locally, i.e. performs more exploitation.
One thing to note from Table 1 is that PSO anecdotically optimises certain airfoils more effectively in both pure and hybrid optimisation scenarios. One possible reason for the behavior in these cases is that if during optimisation with DRL one of the early states of the episode did not converge, making the modifications that the agent does to the optimised airfoil too small. This is, for instance, what happens with NACA63-2a015. Another reason why DRL may perform worse than PSO is because it looks for unrealistic shapes, erring on the side of creativity. This is what appears to happen with airfoil S1223. Finally, there may be airfoils that PSO simply optimises better.
@@
| PSO (hybrid) | 1566 | 64 36 | 150(61) | 0.67 2.21 |
| DRL (hybrid) | 1444 | 89 37 | 179(46) | 11.5 9.4 |
Finally, we provide a comparison of optimisation time per airfoil (averaged over the whole UIUC evaluation set) for DRL and PSO, as performed on an Intel(R) Xeon(R) Gold 6248R CPU @ 3.00GHz+. For concreteness, we focus on the purely aerodynamic case (where the maximum thickness is not necessarily preserved, e.g. ), although the results are qualitatively similar in the hybrid optimisation case. The DRL agent, using 1 process and 1 CPU thread, takes about 23 seconds to optimise all 1566 airfoils, i.e. an average of 0.0147 s/airfoil. Comparatively, PSO was notably slower; actually, too slow to directly perform the optimisation in a single core, so MPI was used to accelerate the evaluation (and then the results per process were inferred). The CPU used has 48 cores, 47 used by worker processes plus the one by master process, and with this setup, optimising all the evaluation set took 22 hours and 26 minutes. Assuming linear acceleration, this means seconds on average per airfoil (using a single process). Accordingly, PSO takes about times more time than DRL to optimise each airfoil using a single process. This is because, once trained, the DRL agent does not need to make any solver call to optimise an airfoil. Since training is a one-time cost, DRL becomes increasingly advantageous when optimising multiple airfoils, as its inference is virtually instantaneous. This highlights the drastic efficiency advantage of DRL, making PSO impractically slow by comparison.
III.2 Hyperparameter exploration in the hybrid optimisation case
Here we provide additional insights on the performance of the DRL agent trained on the hybrid optimisation case (i.e. in Eq. 5). We systematically vary –where larger values of provide more weight to the structural optimisation over the aerodynamic one), and assess (i) the number of training steps needed to reach training convergence, (ii) the lift-to-drag ratio performance metrics improvement and best evaluated on all the airfoils present in the Aerosandbox library and (iii) the variation in the maximum thickness, as a function of the regulariser’s hyperparameter . Results (for the converged airfoils) are depicted in Table 4, which serves as a sensitivity analysis of . It can be observed that the number of necessary training steps varies (typically increases) with . At the same time, results confirm the evidence illustrated in Fig 4 and show that the DRL agent (i) is able to improve the aerodynamic efficiency while (ii) also maintaining the structural integrity of the airfoil. Finally, results also certify that forcing the maximum thickness to be preserved comes at the cost of a smaller improvement of the aerodynamic efficiency.
@@ III.2 Hyperparameter exploration in the hybrid optimisation case
III.3 Enhancing the DRL-based airfoil optimisation with Transfer Learning
To round off, here we compare the training efficiency and evaluate the performance of the trained DRL agent when training is done using a single aerodynamic solver (XFoil alone or, for completeness, NeuralFoil alone) with respect to the case when the agent is pre-trained with NeuralFoil and fine-tuned with XFoil). For the sake of exposition, we separate the purely aerodynamic optimisation ( in Eq. 5) and the hybrid one ().
Purely-aerodynamic optimisation case. Results for all DRL agents in the purely aerodynamic case (with and without transfer learning) are summarised in Table 5, reporting (i) the total number of training steps, (ii) the total solver time, (iii) the number of airfoils evaluated, and the metrics (iv) pointwise statistics of best and (v) improvement over all the airfoils in the evaluation set. Pointwise statistics of best reported in the table are given in terms of median (IQR) as the distributions are not Gaussian (Shapiro-Wilk test) and highly skewed, whereas in the case of improvement we report mean std as the distributions are fairly more symmetrical. Note that to compute the total solver time we simply multiply the number of training steps by the execution time per call to the particular aerodynamic solver, which according to NeuralFoil’s GitHub amounts to ms for a call to XFoil and ms for a call to the smallest version of NeuralFoil, which is the one we actually used. Although the total solver time is not equivalent to the total training time, it is nonetheless an objective way of comparing the computational effort of the agent using vs. not using transfer learning. As previously mentioned, the number of airfoils evaluated may differ slightly between agents because XFoil, which is consistently used for evaluation, occasionally fails to converge. In such cases, we excluded the corresponding airfoils from the analysis.
| DRL Agent | Training steps | Solver time (s) | Airfoils evaluated | Improvement | Best |
| Fully trained with XFoil (no TL) | 81920 | 5980 | 1982 | 241(37) | |
| Pre-training (NeuralFoil) | 26312 | 105 | 1970 | 154(33) | |
| Transfer Learning (strategy #3) | 10240 | 748 | 1983 | 219(29) | |
| Transfer Learning (strategy #1) | 10240 | 748 | 1977 | 236(28) |
First, out of the four transfer learning strategies depicted in Sec. II.3, we found that strategy had the best performance, followed by . Strategies and did not converge and are therefore not reported in the table. This lack of convergence indicates that the information present in the final layers of the neural network plays an important role in shaping the agent’s behaviour.
@@ III.3 Enhancing the DRL-based airfoil optimisation with Transfer Learning
Third, not only the number of steps, but also the total solver time during training ( seconds vs. seconds) is notably shorter for agents trained using transfer learning. On this respect, we have also compared the total training time with and without transfer learning. If we define the Time Reduction () as
we find , reaching higher reductions than in other settings Bhola et al. (2023).
@@
Hybrid optimisation case. Results of the DRL agent (with TL) trained on a combined aerodynamic/structural optimisation are depicted in Table 6. Comparing these with the equivalent results without TL reported in Table 4, we find that (i) the procedure involving TL substantially reduces the number of training steps as well as the training time, reaching Time Reduction as large as ; (ii) fixing the structural hyperparameter , the improvement in aerodynamic efficiency is only slightly lower for every value of (around a 10%) in the TL-enhanced DRL agent than the one without TL, and (iii) the preservation of the maximum thickness is quite similar in both techniques. In summary, TL-enhancement also works in the hybrid optimisation case.